Suppose that the life length (in hours) of a certain radio tube is a continuous random variable Y with p.d.f. f(y) = 100/y2 , 100 < y, zero elsewhere. What is the probability that if 4 such tubes are installed in a set, exactly one will have to be replaced after 150 hours of service?
Answer #1

from above P(a tube last more than 150 hours)=P(Y>150)= f(y) dy = (100/y2) dy =-100/y|150 =100/150 =2/3

hence exactly one will have to be replaced after 150 hours of service =4C1(2/3)1(1/3)3=0.0988